OpenStudy (archtangent):

can i transpose variables in differential equations? x^2 dx + y (x – 1) dy = 0 i want to do this using separation metthod i want to know if it is possible to transpose y(x-1)dy to te other side hence looking lke this after x2 dx + = -y(x+1)dy

ganeshie8 (ganeshie8):

yes thats legit in solving separable eqns

ganeshie8 (ganeshie8):

\[\large x^2 \mathrm{d} x + y (x – 1) \rm{d}y = 0\] \[\large x^2 \mathrm{d} x = y (x – 1) \rm{d}y \] \[\large \dfrac{x^2}{x-1}~ \mathrm{d} x = y ~ \rm{d}y \]

ganeshie8 (ganeshie8):

\[\large \int \dfrac{x^2}{x-1}~ \mathrm{d} x = \int y ~ \rm{d}y \]

OpenStudy (archtangent):

you da real mvp thanks, really helped and appreciate it thanks sir :))

ganeshie8 (ganeshie8):

np :) notice you may need to reduce the degree of numerator using long division for integrating left side integral

OpenStudy (archtangent):

yep, noted thanks sir